A person has four different values of grades in his diploma. Out of these the best value is found twice, the second best is found four times, the second worst is found only once and the worst is found three times. Find the values of these grades, when the mean of the diploma is 7,5 and we know that the values are consecutive whole numbers.


1
Expert's answer
2021-07-13T09:49:13-0400

Let n be the worst value.

"\frac{2(n+3)+4(n+2)+(n+1)+3n}{2+4+1+3}=7.5."

"\frac{10n+15}{10}=7.5."

"10n+15=75."

"n=6."

Values of grades: 6, 7, 8, 9.